OpenStudy (anonymous):

Solve the equation. (cube root of)....2x+13(out of root sign)=-5

OpenStudy (anonymous):

Let me re-type it here.............

OpenStudy (anonymous):

I can't figure out how to type it on here

OpenStudy (anonymous):

is it cuberoot(2x) + 13 = -5

OpenStudy (anonymous):

cuberoot2x+13=-5

OpenStudy (anonymous):

its just cuberoot 2x not cuberoot(2x+13) ?

OpenStudy (anonymous):

It's the cuberoot(2x+13)

OpenStudy (anonymous):

ok cube both sides of equation: 2x + 13 = (-5)^3 = -125 2x = -125-13 = -138 x = -69 thats the solution

OpenStudy (anonymous):

So I just do the third power on the other side?

OpenStudy (anonymous):

\sqrt[3]{x} put it between \[ and

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

nwm, just showing how to type cuberoot

OpenStudy (anonymous):

Oh! Ok, thx!